OpenStudy (anonymous):

Find a value for \(\left\{a\mid0

OpenStudy (anonymous):

um... regardless of what a you choose, b =a will work as a choice for b.

OpenStudy (anonymous):

Should this problem require that b != a?

OpenStudy (anonymous):

Also for an \(a\neq b\). Yup lol forgot

OpenStudy (anonymous):

I think intuitively I can say that numbers like \(\frac{1}{\pi},\frac{1}{e}\) qualify. But I'm not sure how to prove it.

OpenStudy (anonymous):

Honestly, this problem might be a little over my head, because it seems like any a would work. Maybe you can give me an example of an a that does not?

OpenStudy (anonymous):

\[\left(\frac{1}{2}\right)^\frac{1}{2}=\left(\frac{1}{4}\right)^\frac{1}{4}\]

OpenStudy (anonymous):

This is the case for all, I think, non-transcendental numbers since most points are mapped by the function \(x^x\).

OpenStudy (anonymous):

@FoolForMath Halp me :(